### v1

# Mutual benefits of social learning and algorithmic mediation for cumulative culture

Mutual benefits of social learning and algorithmic mediation for cumulative culture
Agnieszka Czaplicka1, Fabian Baumann1,2, Iyad Rahwan1
1Center for Humans and Machines,
Max Planck Institute for Human Development, Lentzeallee 94, Berlin 14195, Germany.
2Department of Biology, University of Pennsylvania, Philadelphia, PA 19104
Corresponding author: agaczapl@gmail.com
The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artefacts that enable us to cope with environmental challenges. The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artefacts (innovation) and copy information from others (social learning). This classic picture overlooks the fact that in the digital age, intelligent algorithms are increasingly mediating information between humans, with potential consequences for cumulative cultural evolution. Building on an established model of cultural evolution, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation. We find that algorithmic mediation has a strong impact on cultural accumulation and that this impact generally increases as social networks become less densely connected. Moreover, cultural accumulation tends to be optimal when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density. Our modeling results are a first step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework. Models of this kind will also help to uncover mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
1 Introduction
Humans’ ability to create complex cultural artefacts such as languages, scientific theories, art and technology, is often seen as one of the most important aspects of our success as a species [1, 2]. An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e. from their peers [3, 4], and pass our their acquired knowledge on to others [3, 5, 1]. Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (prestige bias) [6, 7], to frequent cultural artefacts (context bias) [8, 9] to specific information (content bias) [10]. At the population level, social learning is a necessary condition for cumulative cultural evolution, a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artefacts that lie outside of the scope of individuals [11, 12]. Yet the accumulation of culture in populations can also be limited by factors such as population size [12] or the structure of the underlying social network [13].
In the digital age, much of human social learning is shaped by intelligent algorithms. Instead of learning directly from their peers, individuals absorb information mediated by algorithms. Examples of such algorithms include recommendation systems that influence consumer behavior through personalized product information [14], algorithmically curated content feeds on social media platforms [15] and search engines that rank websites and online content [16]. By facilitating the targeted retrieval of information, these systems can significantly improve the ability of individuals to navigate complex online information landscapes and promote the transfer of knowledge within collectives. However, information-mediating algorithms also have a potential downside: algorithmic personalization has been shown to restrict the scope for individual exploration and limit the diversity of information, for example by actively creating filter bubbles [17] or exacerbating natural tendencies towards homophily [18, 19]. Recognizing these impacts, previous theoretical and empirical works have investigated the impact of algorithms on the diversity of content [20], the formation of opinions [21, 22], and the spread of misinformation [23]. However, little attention has been paid to the question of how algorithmic mediation may affect fundamental and more general processes of cultural evolution [24, 25]. In particular, it is unclear how algorithmic mediation interacts with human social learning to influence the accumulation of complex cultural artifacts.
In this paper, we take a step forward to close this gap by building on an established model of cultural evolution [26]. More specifically, we extend the model by a simple algorithm that first aggregates all available information and then redistributes it in a personalized way. This allows us to study the interplay between social learning (i.e., people share information based on a social network) and algorithmic mediation (i.e., information is mediated by an algorithm). Our modeling results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information transfer. Instead, both types of learning are most powerful when combined, where the optimal ratio depends on the underlying social network. The dynamic interplay between social learning and algorithmic mediation suggests that understanding cultural evolution in the digital age requires considering both types of learning together, rather than in isolation.
2 Model
We build on previous theoretical work [26] that examined how cultural accumulation is affected by different constraints on the individual level. In particular, while the impact of different (social) learning strategies and the costs associated to learning on cultural accumulation were studied, the influence of a social network structure and algorithmic mediation was not investigated.
To fill this gap, we first consider cultural accumulation in populations of agents that communicate based on random (social) networks with different densities of network connections. In a second step, we extend the model by a simple algorithm that mediates information among agents. The individual-based model presented in [26] considers a dynamic process that evolves in generations, with each ”time step” comprising the birth and death of an entire generation of agents. First, the agents copy from the previous generation and then, having learnt all possible information, they start to innovate. Here we deviate from this approach. We modify the model by re-scaling the dynamics so that at each time step a single cultural trait can be acquired either by copying or by innovation [27]. This modification allows us to study the interplay between different mechanisms of information transmission between agents—social learning and algorithmic mediation—and the consequences for population-scale cultural accumulation.
2.1 Cultural space
We consider a space of cultural traits (or cultural space) that defines how new traits (innovations) arise based on existing ones. In line with previous works, the cultural space consists of several cultural branches (or paths), and that traits are functionally dependent, reflecting the cumulative nature of human culture [13, 28, 26, 29].
More specifically, the cultural space is structured as follows. Each trait $x_{l}$ is defined by two components: the cultural branch $x$ it belongs to and its complexity level $l$, see Fig. 1A. Cultural branches $x$ represent the fact that cultural evolution often follows multiple parallel paths of development [28, 13], and levels $l$ model the functional dependency and increasing complexity of cultural artifacts built on top of each other. We consider a finite set of possible branches, i.e., $x=1,...,X$ and assume that cultural complexity levels $l$ are unbounded, i.e. $l=1,2,...,\infty$. The functional dependency of the traits $x$ additionally requires that the traits are acquired in ascending order, starting with the lowest level $l=1$, or more generally: an agent that wants to acquire a trait on level $l$ must possess a trait on level ${l-1}$.
Each cultural trait $x_{l}$ is characterised by an inherent quality score, or payoff, $z(x_{l})$, and we assume that the extent to which the traits are beneficial to agents increases with $z$ [5, 13, 30, 31]. Note, however, that we do not link an individual’s probability to survive to their payoff. Instead, cultural selection happens through selective social learning, where agents copy from agents with high payoffs [30, 32]. Motivated by empirical research on scientific [33] and technological [34] breakthroughs and previous studies on cultural evolution [5], we assume that high-quality (i.e. high-payoff) innovations are rare among many low-quality attempts to innovate. Hence, we draw the payoffs of newly innovated items, $z(x)$, from an exponential distribution, and subsequently square, double, and finally round values to an integer value [27]. This procedure results in a few traits with a payoff of around $z=50$, and roughly half of the traits with $z=0$ [26].
2.2 Dynamics
We consider a population of $N$ agents that are embedded in a social network. Initially, agents are naive, i.e. they do not possess any cultural traits, and their individual effort budget is $B$. A finite effort budget $B$ models the fact that individuals have limited resources, which also constrains cumulative cultural evolution at the collective level. In particular, over time agents spend their budget $B$ on actions to acquire cultural traits, either by copying already discovered traits (learning), or by discovering new traits (innovation). The dynamics of a focal agent $i$ at time $t$ (at development level $l_{i}$ and remaining resources $B(t)$) can be described as a cycle of the following three events:
$(i)$ Death-birth process.
With probability $q=1/N$ agent $i$ is replaced by a naive agent, $i^{\prime}$, which corresponds to the replacement of on average one agent per time step. The initial effort budget of agent $i^{\prime}$ is $B$ and they do not possess any cultural traits. Agent $i^{\prime}$ takes the network position of agent $i$.
$(ii)$ Social learning.
Agent $i$ selects its neighbour $j$ with the highest cumulative payoff [7]. If agent $j$ is at a higher development level than agent $i$, i.e. $l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$-stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget. If the copying of a trait is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e. $l_{j}<l_{i+1}$, the learning costs $C_{s}$ are not deducted from the agent’s budget. Agent $i$ then continues with the innovation step (see below). If the social learning step was successful, the innovation step is skipped. Social learning is schematically depicted in Fig. 1C.
$(iii)$ Innovation.
If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), it will proceed with innovation. First, $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$. If the trait is viable, i.e. its payoff $z(x_{l_{i}+1})>0$, the innovation step is successful and agent $i$ acquires the trait. Note that unsuccessful innovation concludes the time step for agent $i$ without acquiring a new item. Regardless of the success of the innovation step, the cost of innovation $C_{i}$ is subtracted from agent $i$’s effort budget. Innovation is schematically depicted in Fig. 1B.
To quantify the performance of a population with respect to cultural accumulation, we calculate the mean payoff of all agents in the population, i.e. $\bar{Z}(t)=N^{-1}\sum Z_{i}(t)$, where $Z_{i}$ is the cumulative payoff of agent $i$, defined as $Z_{i}=\sum_{l=1}^{l_{\mathrm{max}}}z(x_{l})$. Note that to compute $Z_{i}$, we only sum over the complexity level $l$, as a consequence of the assumed dynamics of innovation and learning: an agent can only acquire a maximum of one trait per complexity level $l$. Additionally, we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait. Furthermore, we follow the intuition that it is easier to copy something from a peer than to innovate it, i.e. we assume $C_{s}<C_{i}$ [26].
2.3 Algorithmic mediation
To investigate how algorithmic mediation can affect cultural accumulation, and to shed light on the interplay between different types of information propagation, i.e. social and algorithm-mediated learning, we extend the model described in the previous section. While still being embedded in a social network, agents can now also learn cultural traits through algorithmic mediation. For simplicity, we implement algorithmic mediation as a ’perfect’ recommender system that has full knowledge about all cultural traits innovated so far in the system, i.e. the cultural reservoirs of every agent. The knowledge about traits’ development levels $(l)$ and their payoffs ($z(x_{l})$) allows the algorithm to provide individuals with personalized recommendations. We assume that learning through algorithmic mediation requires less effort than social learning, i.e., we assume $C_{r}<C_{s}$, where $C_{r}$ defines the costs of algorithmic mediation. In each time step, the agent copies a trait through algorithmic mediation with probability $r$, or otherwise, with probability ($1-r$), the agent learns socially.
Figure 1: Schematic depiction of key model components. A: Open-ended cultural space depicted with four branches ($X=4$), and an infinite number of (complexity) levels $l$. B: Agents innovate a new item on a randomly chosen branch on a level $l+1$. C: Social learning is modeled as a two-step process. (1) Agents choose their neighbor $j$ with the highest cumulative payoff $Z_{j}$. (2) If agent $j$ possesses an item on a sufficient level ($l+1$), the item is copied by agent $i$ and the cost $C_{s}$ is paid. D: Algorithmic mediation is also modeled as a two-step process. (1) The recommender system aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$. (2) Agent $i$ then receives a recommendation of the trait with the highest payoff.
Algorithmic recommendation (with probability $r$).
In every time step, the recommender system aggregates all cultural traits and selects only those on level $l_{i}+1$ (personalization). After sorting those $l_{i}+1$-level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from its budget. Note that when no suitable traits can be recommended, or if agent $i$ does not have enough resources ($B(t)<C_{r}$), the recommendation fails, the agent does not pay a cost, and continues with the innovation step. When the recommendation is successful, agent $i$ does not attempt to innovate. Algorithmic mediation is schematically depicted in Fig. 1D.
Social learning (with probability $1-r$).
Agent $i$ attempts to copy a cultural trait from neighbors (the social learning step is identical to the one described in the previous section).
The order of events in each time step is identical to the one described in the previous section, i.e., consisting of $(i)$ the death-birth process, $(ii)$ learning (either socially or through algorithmic mediation), and $(iii)$ innovation. The different model components are schematically depicted in Fig. 1.
3 Results
If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$ and thus has on average $k=pN$ neighbors. The effort budget of agents is set to $B=1000$, the costs for acquiring cultural traits are fixed at $C_{i}=10$, $C_{s}=5$, and $C_{r}=1$, and we assume $X=100$ cultural branches.
Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the other parameters. In the following, we will thus focus on these ”maximally achievable values”, i.e., population mean values at the stationary state. For brevity, we will refer to these values as ”mean group values” and mark them with a subscript max. Each simulation continues for $t=20N$ time steps and the last $200$ time steps are used to compute the mean group values. Our results are then reported as averages over $200$ realizations.
3.1 Network-based social learning
Figure 2 depicts the results for fully connected networks ($p=1$). Similar to the findings reported in [26], we observe that after the initial growth of $\bar{Z}$ the dynamics reach a quasi-stationary state where the mean group payoff approaches its maximum value, denoted as $\bar{Z}_{\mathrm{max}}$. In line with classical results in cultural evolution [12], we find that the maximum cultural complexity increases with the system size uncovering a logarithmic relationship. This relationship becomes more clear by plotting $\bar{Z}_{\mathrm{max}}$ as a function of $\log(N)$ (system size), as depicted in the right panel of Figure 2.
Figure 2: (left panel) The mean group payoff for fully connected networks as a function of time for different system sizes. (right panel) Maximally achievable group payoff as a function of system size. The remaining model parameters are: $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$.
Figure 3 shows the results for random networks with different connectivity, as parameterized by the average number of neighbours $k$ of an agent. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of cultural traits $\bar{T}_{\mathrm{max}}$, mean group payoff $\bar{Z}_{\mathrm{max}}$. The mean group level grows with the number of neighbors until it reaches a plateau for very dense networks. Instead, the number of discovered items $\bar{T}_{\mathrm{max}}$ decreases rapidly and monotonically with $k$. The most striking behavior is observed for the mean group payoff. Initially, for very sparse networks (low values of $k$) we observe an increase of $\bar{Z}_{\mathrm{max}}$with $k$, but after a maximum value is reached, for intermediate levels of connectivity, we observe that $\bar{Z}_{\mathrm{max}}$ decreases with $k$.
Figure 3: System behavior in a stationary state as a function of network connectivity $k$. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$
3.2 Algorithmic mediation
Next, we focus on the quantitative and qualitative impact of the algorithmic mediation on cumulative cultural evolution. The results are shown in Fig. 4. Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing $r$ for every value of $k$, see left panel of Fig. 4. In line with the observations from the previous section, a higher level of development is observed for denser networks. By contrast, the mean group payoff $\bar{Z}_{\mathrm{max}}$ gives rise to maximum values for intermediate levels of $r$, and these maximum values decrease with $k$. This suggests a trade-off between network connectivity and the ratio of recommendations, which can be seen more clearly in Fig. 5. The highest payoff is observed for very sparse networks (every agent has only one neighbor on average) and relatively high levels of recommendations ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the middle panel of Fig.  4). With fewer neighbors an individual benefits from a higher number of recommendations, but this number decreases with $k$. Note that both limiting cases of $r=0$ and $r=1$ never give rise to maximum values, independently of the connectivity, suggesting that network-based social learning and algorithmic-mediation of information are mutually beneficial.
Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of recommendation probability $r$, and for different network connectivities $k$ including complete graphs (CG). The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and probability of reccomendation $r$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths $X$, do not qualitatively affect our modeling results. We also did not observe non-trivial relationships between the parameters related to learning costs (i.e. $C_{i}$, $C_{s}$, $C_{r}$, $B$) and the measures of cultural accumulation. Intuitively, the more we increase the cost of learning or innovation, the lower the payoff of the group. Similarly, a larger effort budget translates into higher cultural accumulation, where the resulting dynamics depend on the network topology and the recommendation probability $r$. Nevertheless, our findings are applicable within the context of reasonable parameter values, where the costs of acquiring items are significantly lower than the effort budget ($C\ll B$). Finally, it is important to point out that although we focused on cases in which the costs of social learning are higher than the costs of algorithmic mediation, we confirmed that our results are robust for $C_{s}=C_{r}$.
4 Discussion
We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts from the field of cultural evolution [26]. In the first part, we considered the model without algorithmic mediation and found a strong dependence of cultural accumulation on population size. In particular, for fully connected networks, the ability of a population to accumulate culture showed a positive relation (logarithmic growth) with population size, which is consistent with established findings in cultural evolution [6, 35, 36, 37, 38, 39]. We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbors $k$ of a single agent. Although increasing the link density always led to higher levels of development ($\bar{L}_{\mathrm{max}}$), this was not the case for the mean group payoff ($\bar{Z}_{\mathrm{max}}$) and the number of innovated traits ($\bar{T}_{\mathrm{max}}$), which peak at intermediate values of $k$ and monotonously decrease with $k$, respectively. These non-trivial behaviors indicate that cultural accumulation is subject to an exploration-exploitation trade-off between innovation (exploration) and social learning (exploitation), a phenomenon that was previously also observed in both experimental and theoretical works on social dynamics including cultural evolution [13, 28], wisdom of the crowds [40], as well as collective problem-solving [41, 31, 42, 43]. Dense random networks (i.e. networks with high values of $k$) generally allow to disseminate information about cultural traits efficiently (exploitation) and therefore lead to high cultural complexity, however, do so at the expense of cultural diversity (number of traits) as the rate of innovation is impaired due to an over-reliance on social learning. Furthermore, previous empirical and theoretical findings in cultural evolution and social learning studies suggest that collective performance depends on both the complexity of the task and the underlying network structure. In particular, it was found that performance in dense networks is often decreased for complex tasks [31, 42, 43], a phenomenon which is best explained by sufficient levels of transient diversity in less densely connected systems [44]. In Fig. 3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as a function of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks). In other words, populations easily reach high level of development, but do not necessarily do so with high collective payoffs. In our model, collective payoffs are optimal at a medium level of connectivity, where agents optimally solve the explore-exploit trade-off, i.e., have access to the right amount of information through social learning while maintaining sufficient incentives to innovate.
In the second part, we investigated the combined impact of network-based social learning and algorithmic mediation on cultural accumulation. For a fixed value of $k$, we find that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the developmental level monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$. The less we are connected, the more we can benefit from algorithmic mediation. For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much of effective connectivity which decreases performance as $r\rightarrow 1$. This is in line, with the strongly decreasing number of traits as $r$ increases, indicating low levels of innovation. These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g. payoff-biased learning). In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and lesser cultural traits than a combination of both processes. This behavior may point toward a general trade-off in socio-technical systems between access to information (and thus the possibility to exchange it) and the ability of individuals to innovate: while algorithms can be beneficial for cultural accumulation, they may also lead to detrimental effects if used excessively [45, 46].
Our research takes a first step in modeling the impact of personalized algorithmic mediation on cumulative cultural evolution, however, it does not come without limitations. By focusing on the algorithmic mediation of information, inspired by recommender systems or feed algorithms, we have not considered other types of intelligent algorithms that are increasingly intertwined with human culture, in particular, generative artificial intelligence [47]. Importantly, generative AI not only re-distributes information, but is able to create novel cultural variation, which is not captured in our models. Furthermore, in line with previous models, we considered a very simplistic cultural space, namely cultural branches that are independent from each other. A more realistic approach could therefore take into account that cultural innovations often arise through the recombination of artefacts from distant domains, i.e. modeling interdependencies between branches. Similarly, by investigating random networks and agents with identical social learning strategies we have omitted many of the behavioral and structural heterogeneities observed in real social systems [48, 49]. The use of more complex network models—such as scale-free [50] and small-world networks [51]—as well as agents with individual preferences are promising avenues for future research directions. Finally, we have operationalised algorithmic mediation as a simple system that provides agents with ”ideal” personalised recommendations. Therefore, extending our approach to more realistic models of algorithmic mediation, e.g. recommendations based on collaborative filtering [52], is a promising next step. Overall, our work is only a first step towards a better theoretical understanding of the impact of algorithms on cumulative cultural evolution. Ultimately, the predictions of our model must stand up to empirical testing, both through controlled laboratory experiments and real-world observational studies across different online environments.
Code availability
Simulation code is available at https://github.com/aga-cz/Mutual-benefits-of-SL-and-RS

### v2

@@
 Mutual benefits of social learning and algorithmic mediation for cumulative culture
-Agnieszka Czaplicka1, Fabian Baumann1,2, Iyad Rahwan1
+Agnieszka Czaplicka1,2,∗, Fabian Baumann1,3, Iyad Rahwan1
 1Center for Humans and Machines,
 Max Planck Institute for Human Development, Lentzeallee 94, Berlin 14195, Germany.
-2Department of Biology, University of Pennsylvania, Philadelphia, PA 19104
+2Faculty of Physics, Warsaw University of Technology, Warszawa, Poland.
+3Department of Biology, University of Pennsylvania, Philadelphia, PA 19104.
 Corresponding author: agaczapl@gmail.com
-The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artefacts that enable us to cope with environmental challenges.
-The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artefacts (innovation) and copy information from others (social learning).
-This classic picture overlooks the fact that in the digital age, intelligent algorithms are increasingly mediating information between humans, with potential consequences for cumulative cultural evolution.
-Building on an established model of cultural evolution, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation.
-We find that algorithmic mediation has a strong impact on cultural accumulation and that this impact generally increases as social networks become less densely connected.
-Moreover, cultural accumulation tends to be optimal when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density.
-Our modeling results are a first step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework.
-Models of this kind will also help to uncover mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
+The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artifacts that enable us to cope with environmental challenges.
+The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artifacts (innovation) and copy information from others (social learning).
+This classic picture overlooks the growing role of intelligent algorithms in the digital age (e.g., search engines, recommender systems, large language models) in mediating information between humans, with potential consequences for cumulative cultural evolution.
+Building on a previous model, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation.
+We find that algorithmic mediation significantly impacts cultural accumulation and that this impact grows as social networks become less densely connected.
+Cultural accumulation is most effective when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density.
+This work is an initial step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework.
+Models like ours provide insights into mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
 1 Introduction
-Humans’ ability to create complex cultural artefacts such as languages, scientific theories, art and technology, is often seen as one of the most important aspects of our success as a species [1, 2].
-An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e.
-from their peers [3, 4], and pass our their acquired knowledge on to others [3, 5, 1].
-Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (prestige bias) [6, 7], to frequent cultural artefacts (context bias) [8, 9] to specific information (content bias) [10].
-At the population level, social learning is a necessary condition for cumulative cultural evolution, a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artefacts that lie outside of the scope of individuals [11, 12].
-Yet the accumulation of culture in populations can also be limited by factors such as population size [12] or the structure of the underlying social network [13].
-In the digital age, much of human social learning is shaped by intelligent algorithms.
-Instead of learning directly from their peers, individuals absorb information mediated by algorithms.
-Examples of such algorithms include recommendation systems that influence consumer behavior through personalized product information [14], algorithmically curated content feeds on social media platforms [15] and search engines that rank websites and online content [16].
-By facilitating the targeted retrieval of information, these systems can significantly improve the ability of individuals to navigate complex online information landscapes and promote the transfer of knowledge within collectives.
-However, information-mediating algorithms also have a potential downside: algorithmic personalization has been shown to restrict the scope for individual exploration and limit the diversity of information, for example by actively creating filter bubbles [17] or exacerbating natural tendencies towards homophily [18, 19].
-Recognizing these impacts, previous theoretical and empirical works have investigated the impact of algorithms on the diversity of content [20], the formation of opinions [21, 22], and the spread of misinformation [23].
-However, little attention has been paid to the question of how algorithmic mediation may affect fundamental and more general processes of cultural evolution [24, 25].
-In particular, it is unclear how algorithmic mediation interacts with human social learning to influence the accumulation of complex cultural artifacts.
-In this paper, we take a step forward to close this gap by building on an established model of cultural evolution [26].
-More specifically, we extend the model by a simple algorithm that first aggregates all available information and then redistributes it in a personalized way.
-This allows us to study the interplay between social learning (i.e., people share information based on a social network) and algorithmic mediation (i.e., information is mediated by an algorithm).
-Our modeling results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information transfer.
+Humans’ ability to create complex cultural artifacts such as languages, scientific theories, art, and technology (including physical artifacts) is often seen as one of the most important aspects of our success as a species [1, 2].
+An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e., copy information from their peers [1, 3, 4, 5].
+Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (payoff bias) [6, 7, 8], to frequent cultural artifacts (context bias) [9, 10], to specific information (content bias) [11].
+At the population level, social learning is a necessary condition for cumulative cultural evolution (or cultural accumulation), a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artifacts that lie outside of the scope of individuals [12, 13].
+Yet the cultural accumulation can also be limited by factors such as population size [13], the structure of the underlying social network [14], or the social learning strategies that humans use [15].
+In the digital age, cultural accumulation is increasingly shaped by algorithmic mediation [16].
+For the purposes of this paper, we broadly define algorithmic mediation as a collection of processes through which intelligent systems aggregate and redistribute information on an unprecedented scale, thereby impacting cultural accumulation.
+These algorithms include recommender systems that influence consumer behavior [17], content-ranking mechanisms on social media platforms [18], and search engines like Google [19].
+More recently, large language models (LLMs) (e.g., GPT, BERT, Gemini, and others) have emerged as transformative tools, reshaping how we acquire knowledge and develop skills.
+Unlike earlier recommender systems, which primarily tackled e-commerce challenges by focusing on consumer preferences, LLMs are likely influencing a far broader range of domains that are central to cultural accumulation.
+These include generating ideas [20], facilitating problem-solving [21], assisting with technical tasks such as coding [22], and offering personalized explanations in fields like mathematics and science [23].
+By synthesizing vast repositories of information and tailoring information to individual needs, algorithmic systems have indeed the potential to impact cumulative cultural evolution by enhancing innovation rates across domains, including technology and basic science [24], reinforcing filter bubbles [25], and limiting the diversity of information shared [26].
+Recognizing these impacts, previous theoretical and empirical works on socio-technological systems have mainly focused on the impact of algorithms on the diversity of content [27], the formation of opinions [28, 29], and the spread of misinformation [30].
+However, the question of how algorithmic mediation may affect cultural accumulation received much less attention.
+To close this gap, we build on—and extend—an established model of cultural evolution [31], where algorithmic mediation is conceptualized as a mechanism that aggregates and redistributes cultural traits on a collective level.
+By integrating social learning and algorithmic mediation in a single modeling framework, we are able to capture the dynamics induced by the interaction of localized peer influence (social learning) and global information redistribution (algorithmic mediation).
+Our results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information.
 Instead, both types of learning are most powerful when combined, where the optimal ratio depends on the underlying social network.
-The dynamic interplay between social learning and algorithmic mediation suggests that understanding cultural evolution in the digital age requires considering both types of learning together, rather than in isolation.
+This insight underscores the need to examine algorithmic mediation and social learning as interconnected forces shaping cultural evolution in the digital age, rather than as isolated processes.
 2 Model
-We build on previous theoretical work [26] that examined how cultural accumulation is affected by different constraints on the individual level.
-In particular, while the impact of different (social) learning strategies and the costs associated to learning on cultural accumulation were studied, the influence of a social network structure and algorithmic mediation was not investigated.
-To fill this gap, we first consider cultural accumulation in populations of agents that communicate based on random (social) networks with different densities of network connections.
-In a second step, we extend the model by a simple algorithm that mediates information among agents.
-The individual-based model presented in [26] considers a dynamic process that evolves in generations, with each ”time step” comprising the birth and death of an entire generation of agents.
-First, the agents copy from the previous generation and then, having learnt all possible information, they start to innovate.
-Here we deviate from this approach.
-We modify the model by re-scaling the dynamics so that at each time step a single cultural trait can be acquired either by copying or by innovation [27].
-This modification allows us to study the interplay between different mechanisms of information transmission between agents—social learning and algorithmic mediation—and the consequences for population-scale cultural accumulation.
+We build on prior theoretical work that explored how cultural accumulation is influenced by various aspects of individual learning, including social learning strategies and the costs associated with learning [31].
+First, we extend this model by incorporating an explicit social network structure.
+In a second step, we extend the model by a simple algorithm that mediates information exchange among agents.
+The individual-based model presented in [31] simulates a dynamic process that unfolds over generations, where each ”time step” represents the simultaneous birth and death of an entire generation of $N$ agents.
+In this framework, agents first copy information from the previous generation and then, after acquiring all available knowledge, begin to innovate.
+In this work, we deviate from this approach and modify the model by re-scaling the dynamics: at each time step, a single cultural trait is acquired either through copying or innovation [32].
+This modification enables us to examine the interplay between different mechanisms of information transmission—social learning and algorithmic mediation—and their implications for population-scale cultural accumulation.
+Below, we describe our modeling approach and provide additional implementation details in the Supplemental Material (SM).
 2.1 Cultural space
 We consider a space of cultural traits (or cultural space) that defines how new traits (innovations) arise based on existing ones.
-In line with previous works, the cultural space consists of several cultural branches (or paths), and that traits are functionally dependent, reflecting the cumulative nature of human culture [13, 28, 26, 29].
-More specifically, the cultural space is structured as follows.
+In line with previous works, the cultural space consists of several cultural branches (or paths), and traits are functionally dependent, reflecting the cumulative nature of human culture [14, 33, 31, 34].
+The cultural space is structured as follows.
 Each trait $x_{l}$ is defined by two components: the cultural branch $x$ it belongs to and its complexity level $l$, see Fig.
 1A.
-Cultural branches $x$ represent the fact that cultural evolution often follows multiple parallel paths of development [28, 13], and levels $l$ model the functional dependency and increasing complexity of cultural artifacts built on top of each other.
+Cultural branches $x$ model the fact that cultural evolution often follows multiple parallel paths [33, 14], and levels $l$ model the functional dependency and increasing complexity of cultural traits built on top of each other.
 We consider a finite set of possible branches, i.e., $x=1,...,X$ and assume that cultural complexity levels $l$ are unbounded, i.e.
-$l=1,2,...,\infty$.
+$l=1,2,...,\infty$, i.e.
+cultural artifacts can be modified infinite number of times.
 The functional dependency of the traits $x$ additionally requires that the traits are acquired in ascending order, starting with the lowest level $l=1$, or more generally: an agent that wants to acquire a trait on level $l$ must possess a trait on level ${l-1}$.
@@
 This procedure results in a few traits with a payoff of around $z=50$, and roughly half of the traits with $z=0$ [31].
+The distinction between complexity levels $(l)$ and payoffs $(z)$ in the model captures the relationship between the accumulation of culture (i.e.
+knowledge, skills, or innovations) and its practical utility in real-world cultural dynamics.
+For instance, consider the development of a new scientific method (skill) or technological tool (physical artifact).
+The complexity of a tool—reflected in its levels—might represent the number of iterative modifications and refinements over time, such as the progression from early calculators to modern computers.
+However, not all refinements enhance utility or payoff.
+Some modifications, like aesthetic improvements or added features, may increase the complexity of a tool without necessarily maximizing its functional effectiveness or societal impact.
+Similarly, in scientific endeavors, researchers may pursue increasingly complex methodologies that yield incremental theoretical insights, but only some of these contribute substantially to practical applications or yield high-impact results capable of shifting the paradigm within a discipline.
+This distinction between levels and payoffs underscores the importance to disentangle how cumulative cultural processes lead to either optimal or suboptimal outcomes, depending on the social and algorithmic mechanisms mediating cultural transmission.
 2.2 Dynamics
@@
 A finite effort budget $B$ models the fact that individuals have limited resources, which also constrains cumulative cultural evolution at the collective level.
-In particular, over time agents spend their budget $B$ on actions to acquire cultural traits, either by copying already discovered traits (learning), or by discovering new traits (innovation).
-The dynamics of a focal agent $i$ at time $t$ (at development level $l_{i}$ and remaining resources $B(t)$) can be described as a cycle of the following three events:
+Over time, agents acquire cultural traits either by copying existing traits (learning) or by discovering new ones (innovation), with each action incurring a cost.
+The dynamics of a focal agent $i$ at time $t$, characterized by the highest complexity level $l_{i}$ of their traits and their remaining resources $B(t)$, can be modeled as a cycle consisting of the following three events:
 $(i)$ Death-birth process.
@@
 $(ii)$ Social learning.
-Agent $i$ selects its neighbour $j$ with the highest cumulative payoff [7].
-If agent $j$ is at a higher development level than agent $i$, i.e.
-$l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$-stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget.
-If the copying of a trait is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e.
-$l_{j}<l_{i+1}$, the learning costs $C_{s}$ are not deducted from the agent’s budget.
+Agent $i$ selects their neighbour $j$ with the highest cumulative payoff [7].
+If agent $j$ is at a higher complexity level $l$ than agent $i$, i.e.
+$l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$–stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget.
+If copying is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e.
+$l_{j}<l_{i+1}$, the learning cost $C_{s}$ is not deducted from the agent’s budget.
 Agent $i$ then continues with the innovation step (see below).
@@
 1C.
+We mainly focus on payoff-biased learning as described above, however, in the SM we also show results for other social learning strategies.
 $(iii)$ Innovation.
-If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), it will proceed with innovation.
-First, $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$.
+If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), they will proceed with innovation.
+First, agent $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$.
 If the trait is viable, i.e.
@@
 1B.
-To quantify the performance of a population with respect to cultural accumulation, we calculate the mean payoff of all agents in the population, i.e.
-$\bar{Z}(t)=N^{-1}\sum Z_{i}(t)$, where $Z_{i}$ is the cumulative payoff of agent $i$, defined as $Z_{i}=\sum_{l=1}^{l_{\mathrm{max}}}z(x_{l})$.
-Note that to compute $Z_{i}$, we only sum over the complexity level $l$, as a consequence of the assumed dynamics of innovation and learning: an agent can only acquire a maximum of one trait per complexity level $l$.
-Additionally, we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait.
+To analyze cultural accumulation in the model, we use three distinct measures.
+First, we calculate the mean payoff of all agents in the population, defined as $\bar{Z}(t)=\frac{1}{N}\sum\limits_{i}Z_{i}(t)$, where $Z_{i}$ represents the cumulative payoff of agent $i$, given by $Z_{i}=\sum\limits_{l=1}^{l_{\mathrm{max}}}z(x_{l})$.
+Here, $z(x_{l})$ denotes the payoff associated with trait $x_{l}$, and the summation is over the complexity levels $l$, reflecting the model’s dynamics: each agent can acquire at most one trait per complexity level.
+The mean payoff $\bar{Z}(t)$ serves as a proxy for the performance of the population, indicating the overall utility derived from the accumulated traits.
+Second, we count the total number of distinct traits created during the process.
+This measure provides an indicator of the diversity of traits in cultural accumulation.
+Finally, we track the highest levels of traits reached by agents, which assess the average cultural complexity within the population.
+These three measures—performance, diversity, and complexity—together offer a comprehensive framework for analyzing the dynamics of cultural accumulation in the model.
+Note that we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait.
 Furthermore, we follow the intuition that it is easier to copy something from a peer than to innovate it, i.e.
@@
 2.3 Algorithmic mediation
-To investigate how algorithmic mediation can affect cultural accumulation, and to shed light on the interplay between different types of information propagation, i.e.
-social and algorithm-mediated learning, we extend the model described in the previous section.
-While still being embedded in a social network, agents can now also learn cultural traits through algorithmic mediation.
-For simplicity, we implement algorithmic mediation as a ’perfect’ recommender system that has full knowledge about all cultural traits innovated so far in the system, i.e.
-the cultural reservoirs of every agent.
-The knowledge about traits’ development levels $(l)$ and their payoffs ($z(x_{l})$) allows the algorithm to provide individuals with personalized recommendations.
-We assume that learning through algorithmic mediation requires less effort than social learning, i.e., we assume $C_{r}<C_{s}$, where $C_{r}$ defines the costs of algorithmic mediation.
-In each time step, the agent copies a trait through algorithmic mediation with probability $r$, or otherwise, with probability ($1-r$), the agent learns socially.
+To investigate how algorithmic mediation affects cultural accumulation and the interplay between social and algorithm-mediated learning, we extend the model described in the previous section.
+While still embedded in a social network, agents can now also learn cultural traits through algorithmic mediation.
+We implement this as a ’perfect’ algorithm with full knowledge of all cultural traits innovated so far, including their development levels $(l)$ and payoffs ($z$), enabling personalization of information.
+Learning through algorithmic mediation is assumed to be less costly than social learning ($C_{r}<C_{s}$).
+We implement the balance between social and algorithmic learning as a stochastic process: At each time step, an agent copies a trait via algorithmic mediation with probability $r$ or learns socially with probability $(1-r)$.
 Figure 1: Schematic depiction of key model components.
@@
 C: Social learning is modeled as a two-step process.
-(1) Agents choose their neighbor $j$ with the highest cumulative payoff $Z_{j}$.
+(1) Agents choose their neighbour $j$ with the highest cumulative payoff $Z_{j}$.
 (2) If agent $j$ possesses an item on a sufficient level ($l+1$), the item is copied by agent $i$ and the cost $C_{s}$ is paid.
 D: Algorithmic mediation is also modeled as a two-step process.
-(1) The recommender system aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$.
+(1) The algorithm aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$.
 (2) Agent $i$ then receives a recommendation of the trait with the highest payoff.
 Algorithmic recommendation (with probability $r$).
-In every time step, the recommender system aggregates all cultural traits and selects only those on level $l_{i}+1$ (personalization).
-After sorting those $l_{i}+1$-level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from its budget.
+In every time step, the algorithm aggregates all cultural traits and selects only those on level $l_{i}\!+\!1$ (personalization).
+After sorting those $l_{i}\!+\!1$–level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from their budget.
 Note that when no suitable traits can be recommended, or if agent $i$ does not have enough resources ($B(t)<C_{r}$), the recommendation fails, the agent does not pay a cost, and continues with the innovation step.
@@
 Social learning (with probability $1-r$).
-Agent $i$ attempts to copy a cultural trait from neighbors (the social learning step is identical to the one described in the previous section).
+Agent $i$ attempts to copy a cultural trait from their neighbours, as described in the previous section.
 The order of events in each time step is identical to the one described in the previous section, i.e., consisting of $(i)$ the death-birth process, $(ii)$ learning (either socially or through algorithmic mediation), and $(iii)$ innovation.
@@
 1.
+Details of the implementation are presented in the SM.
+While idealized, our model of algorithmic mediation broadly mirrors the capabilities of advanced algorithmic systems, such as large language models (LLMs), which tailor information delivery to align with user needs in order to maximize utility, although their effectiveness remains constrained by current methodological and technological limitations.
+For instance, reinforcement learning from human feedback, as employed in systems like ChatGPT, enables the prioritization of useful, accurate, and contextually relevant information.
+While this process is not directly analogous to biological fitness proxies in traditional cultural evolution models, it effectively captures the adaptive refinement and redistribution of cultural traits—including knowledge and skills—within an algorithmically mediated environment.
 3 Results
-If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$ and thus has on average $k=pN$ neighbors.
+If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$.
+Hence, each agent has on average $k=p(N-1)$ neighbours.
 The effort budget of agents is set to $B=1000$, the costs for acquiring cultural traits are fixed at $C_{i}=10$, $C_{s}=5$, and $C_{r}=1$, and we assume $X=100$ cultural branches.
-Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the other parameters.
-In the following, we will thus focus on these ”maximally achievable values”, i.e., population mean values at the stationary state.
+The SM contains a sensitivity analysis of the model with respect to these parameters.
+Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the remaining parameters.
+In the following, we will thus focus on ”maximally achievable values”, i.e., population mean values at the stationary state for performance ($\bar{Z}$), diversity ($\bar{T}$), and complexity ($\bar{L}$).
 For brevity, we will refer to these values as ”mean group values” and mark them with a subscript max.
 Each simulation continues for $t=20N$ time steps and the last $200$ time steps are used to compute the mean group values.
-Our results are then reported as averages over $200$ realizations.
+Our results are reported as averages over $200$ realizations.
 3.1 Network-based social learning
@@
 Similar to the findings reported in [31], we observe that after the initial growth of $\bar{Z}$ the dynamics reach a quasi-stationary state where the mean group payoff approaches its maximum value, denoted as $\bar{Z}_{\mathrm{max}}$.
-In line with classical results in cultural evolution [12], we find that the maximum cultural complexity increases with the system size uncovering a logarithmic relationship.
+In line with classical results in cultural evolution [13], we find that the maximum group performance increases with the system size uncovering a logarithmic relationship.
 This relationship becomes more clear by plotting $\bar{Z}_{\mathrm{max}}$ as a function of $\log(N)$ (system size), as depicted in the right panel of Figure 2.
@@
 Figure 3 shows the results for random networks with different connectivity, as parameterized by the average number of neighbours $k$ of an agent.
-From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of cultural traits $\bar{T}_{\mathrm{max}}$, mean group payoff $\bar{Z}_{\mathrm{max}}$.
-The mean group level grows with the number of neighbors until it reaches a plateau for very dense networks.
+From left to right we depict the mean group level $\bar{L}_{\mathrm{max}}$, the mean number of cultural traits $\bar{T}_{\mathrm{max}}$, and the mean group payoff $\bar{Z}_{\mathrm{max}}$ as functions of $k$.
+The mean group level grows with the number of neighbours until it reaches a plateau for very dense networks.
 Instead, the number of discovered items $\bar{T}_{\mathrm{max}}$ decreases rapidly and monotonically with $k$.
@@
 Initially, for very sparse networks (low values of $k$) we observe an increase of $\bar{Z}_{\mathrm{max}}$with $k$, but after a maximum value is reached, for intermediate levels of connectivity, we observe that $\bar{Z}_{\mathrm{max}}$ decreases with $k$.
-Figure 3: System behavior in a stationary state as a function of network connectivity $k$.
+Figure 3: Stationary states as functions of network connectivity $k$.
 From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$.
@@
 3.2 Algorithmic mediation
-Next, we focus on the quantitative and qualitative impact of the algorithmic mediation on cumulative cultural evolution.
-The results are shown in Fig.
-4.
-Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing $r$ for every value of $k$, see left panel of Fig.
+In Fig.
+4, we show the results for the impact of algorithmic mediation on cumulative cultural evolution.
+Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing ratio of algorithmic mediation $r$ for every value of $k$, see left panel of Fig.
 4.
@@
 By contrast, the mean group payoff $\bar{Z}_{\mathrm{max}}$ gives rise to maximum values for intermediate levels of $r$, and these maximum values decrease with $k$.
-This suggests a trade-off between network connectivity and the ratio of recommendations, which can be seen more clearly in Fig.
+This suggests a trade-off between the network connectivity ($k$) and the ratio of algorithmic mediation ($r$), which can be seen more clearly in Fig.
 5.
-The highest payoff is observed for very sparse networks (every agent has only one neighbor on average) and relatively high levels of recommendations ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the middle panel of Fig.
-4).
-With fewer neighbors an individual benefits from a higher number of recommendations, but this number decreases with $k$.
+The highest payoff is observed for very sparse networks (every agent has only one neighbour on average) and relatively high levels of algorithmic mediation ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the bottom panel of Fig.
+4.
+With fewer neighbours an individual benefits from a higher level of algorithmic mediation, but this trend decreases with $k$.
 Note that both limiting cases of $r=0$ and $r=1$ never give rise to maximum values, independently of the connectivity, suggesting that network-based social learning and algorithmic-mediation of information are mutually beneficial.
-Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of recommendation probability $r$, and for different network connectivities $k$ including complete graphs (CG).
+Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of the ratio of algorithmic mediation $r$, and for different network connectivities $k$ including complete graphs (CG).
 The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
-Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and probability of reccomendation $r$.
-The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
-From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths $X$, do not qualitatively affect our modeling results.
-We also did not observe non-trivial relationships between the parameters related to learning costs (i.e.
-$C_{i}$, $C_{s}$, $C_{r}$, $B$) and the measures of cultural accumulation.
-Intuitively, the more we increase the cost of learning or innovation, the lower the payoff of the group.
-Similarly, a larger effort budget translates into higher cultural accumulation, where the resulting dynamics depend on the network topology and the recommendation probability $r$.
-Nevertheless, our findings are applicable within the context of reasonable parameter values, where the costs of acquiring items are significantly lower than the effort budget ($C\ll B$).
-Finally, it is important to point out that although we focused on cases in which the costs of social learning are higher than the costs of algorithmic mediation, we confirmed that our results are robust for $C_{s}=C_{r}$.
+Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and the ratio of algorithmic mediation $r$.
+The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$.
+From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths ($X$), do not qualitatively affect our modeling results.
+We also did not observe non-trivial relationships between the effort budget ($B$) and the measures of cultural accumulation.
+Intuitively, a larger effort budget translates into higher payoffs, where the resulting dynamics depend on the network topology and the ratio of algorithmic mediation $r$.
+In the SM, we report additional simulation results on how the costs of acquiring cultural traits (i.e.
+$C_{i}$, $C_{s}$, $C_{r}$) affect the dynamics.
+We find that lower values of the learning costs are associated with higher measures of cultural accumulation, where the impact of ($C_{i}$, $C_{s}$, $C_{r}$) is most pronounced for the mean group level, and much smaller for the mean number of traits and the mean group payoff.
+We also noticed some qualitative changes of $\bar{Z}_{max}$ as a function of the ratio of algorithmic mediation $r$ for different values of $C_{i}$, $C_{s}$, and $C_{r}$.
+Importantly, we find that in every investigated case our main result remains unchanged, i.e.
+intermediate values of $r$ give rise to the maximum performance (peak values of $\bar{Z}_{max}$), where the exact location of that optimum depends on the costs of innovation, social learning and algorithmic mediation, as reported in the SM.
 4 Discussion
-We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts from the field of cultural evolution [26].
+We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts in cultural evolution [31].
 In the first part, we considered the model without algorithmic mediation and found a strong dependence of cultural accumulation on population size.
-In particular, for fully connected networks, the ability of a population to accumulate culture showed a positive relation (logarithmic growth) with population size, which is consistent with established findings in cultural evolution [6, 35, 36, 37, 38, 39].
-We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbors $k$ of a single agent.
+In particular, for fully connected networks, the population’s ability to accumulate culture exhibited a positive relationship with population size, characterized by logarithmic growth.
+This result aligns with established findings in cultural evolution, highlighting the role of population size in fostering cumulative cultural dynamics [6, 40, 41, 42, 43, 44].
+We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbours $k$ of a single agent.
 Although increasing the link density always led to higher levels of development ($\bar{L}_{\mathrm{max}}$), this was not the case for the mean group payoff ($\bar{Z}_{\mathrm{max}}$) and the number of innovated traits ($\bar{T}_{\mathrm{max}}$), which peak at intermediate values of $k$ and monotonously decrease with $k$, respectively.
 These non-trivial behaviors indicate that cultural accumulation is subject to an exploration-exploitation trade-off between innovation (exploration) and social learning (exploitation), a phenomenon that was previously also observed in both experimental and theoretical works on social dynamics including cultural evolution [14, 33], wisdom of the crowds [45], as well as collective problem-solving [46, 36, 47, 48].
-Dense random networks (i.e.
-networks with high values of $k$) generally allow to disseminate information about cultural traits efficiently (exploitation) and therefore lead to high cultural complexity, however, do so at the expense of cultural diversity (number of traits) as the rate of innovation is impaired due to an over-reliance on social learning.
+Dense random networks (i.e., networks with high values of $k$) generally allow for the efficient dissemination of information about cultural traits (exploitation) and therefore lead to high cultural complexity.
+However, this comes at the expense of cultural diversity (number of traits), as the rate of innovation is impaired due to an over-reliance on social learning.
 Furthermore, previous empirical and theoretical findings in cultural evolution and social learning studies suggest that collective performance depends on both the complexity of the task and the underlying network structure.
-In particular, it was found that performance in dense networks is often decreased for complex tasks [31, 42, 43], a phenomenon which is best explained by sufficient levels of transient diversity in less densely connected systems [44].
+In particular, it was found that performance in dense networks is often decreased for complex tasks [36, 47, 48], a phenomenon best explained by sufficient levels of transient diversity in less densely connected systems [49].
 In Fig.
-3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as a function of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks).
-In other words, populations easily reach high level of development, but do not necessarily do so with high collective payoffs.
+3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as functions of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks).
+In other words, populations develop traits and reach high complexity rather easily, however, such traits do not necessarily have a high utility (i.e.
+high payoff).
 In our model, collective payoffs are optimal at a medium level of connectivity, where agents optimally solve the explore-exploit trade-off, i.e., have access to the right amount of information through social learning while maintaining sufficient incentives to innovate.
 In the second part, we investigated the combined impact of network-based social learning and algorithmic mediation on cultural accumulation.
-For a fixed value of $k$, we find that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the developmental level monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$.
+For a fixed value of $k$, we found that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the cultural complexity monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$.
 The less we are connected, the more we can benefit from algorithmic mediation.
-For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much of effective connectivity which decreases performance as $r\rightarrow 1$.
-This is in line, with the strongly decreasing number of traits as $r$ increases, indicating low levels of innovation.
-These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g.
-payoff-biased learning).
-In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and lesser cultural traits than a combination of both processes.
+For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much effective connectivity, which decreases performance as $r\rightarrow 1$.
+This result is in line with the strongly decreasing number of traits as $r$ increases, indicating low levels of cultural diversity.
+Interestingly, this behavior remains remarkably stable when the structure of the underlying network is altered, provided the link density is kept constant, as demonstrated in the Supplemental Material (SM).
+For two additional network structures—small-world networks [50] and scale-free networks [51]—the mean group level, the number of traits, and the payoffs exhibit the same qualitative behavior as observed in random ER networks.
+Furthermore, we examined alternative social learning strategies (in addition to the payoff-biased learning addressed previously) and their influence on cultural accumulation in the presence of algorithmic mediation.
+Two strategies were considered: one in which cultural traits are copied from the neighbour with the highest complexity level, and alternatively, one in which a random neighbour is chosen.
+The latter strategy is analogous to the unbiased transmission described in [31].
+It can be seen that for $r=0$ (solely social learning), the random strategy performs considerably less well than the other two, but the situation becomes much more intricate when algorithmic mediation is introduced (see Figure 2 in SM).
+The mean number of levels as a function of $r$ is considerably lower for the random selection strategy, yet it yields larger payoffs for high values of $r$.
+This seemingly contradictory effect is the result of two mechanisms: firstly, a randomly selected neighbour is rarely at a sufficiently high level to allow for copying, which leads to a high number of innovations (as evidenced by the values of the mean number of traits in the middle panel of Figure 2 in SM); secondly, when the proportion of algorithmic recommendations is sufficiently high, the innovated items are efficiently distributed throughout the system.
+Consequently, a seemingly not effective random strategy at the individual level becomes optimal for the population.
+These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g., payoff-biased learning).
+In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and fewer cultural traits than a combination of both processes.
 This behavior may point toward a general trade-off in socio-technical systems between access to information (and thus the possibility to exchange it) and the ability of individuals to innovate: while algorithms can be beneficial for cultural accumulation, they may also lead to detrimental effects if used excessively [52, 53].
-Our research takes a first step in modeling the impact of personalized algorithmic mediation on cumulative cultural evolution, however, it does not come without limitations.
-By focusing on the algorithmic mediation of information, inspired by recommender systems or feed algorithms, we have not considered other types of intelligent algorithms that are increasingly intertwined with human culture, in particular, generative artificial intelligence [47].
-Importantly, generative AI not only re-distributes information, but is able to create novel cultural variation, which is not captured in our models.
-Furthermore, in line with previous models, we considered a very simplistic cultural space, namely cultural branches that are independent from each other.
-A more realistic approach could therefore take into account that cultural innovations often arise through the recombination of artefacts from distant domains, i.e.
-modeling interdependencies between branches.
-Similarly, by investigating random networks and agents with identical social learning strategies we have omitted many of the behavioral and structural heterogeneities observed in real social systems [48, 49].
-The use of more complex network models—such as scale-free [50] and small-world networks [51]—as well as agents with individual preferences are promising avenues for future research directions.
-Finally, we have operationalised algorithmic mediation as a simple system that provides agents with ”ideal” personalised recommendations.
-Therefore, extending our approach to more realistic models of algorithmic mediation, e.g.
-recommendations based on collaborative filtering [52], is a promising next step.
-Overall, our work is only a first step towards a better theoretical understanding of the impact of algorithms on cumulative cultural evolution.
-Ultimately, the predictions of our model must stand up to empirical testing, both through controlled laboratory experiments and real-world observational studies across different online environments.
-Code availability
-Simulation code is available at https://github.com/aga-cz/Mutual-benefits-of-SL-and-RS
+The real-world counterparts of these dynamics provide valuable insights into modern socio-technical systems.
+For example, recommender systems and large language models (LLMs) shape cultural accumulation by redistributing information globally, facilitating knowledge acquisition and skill development.
+However, they also risk suppressing diversity or incentivizing conformity when over-relied upon.
+Applications in collaborative platforms [54, 55], professional networks [56], or online communities [30, 57] could serve as real-world analogs of the processes modeled here.
+Empirical studies might examine how varying levels of algorithmic support in such environments affect innovation rates or cultural diversity, shedding light on the interplay between cultural exploration and exploitation in digital ecosystems [58, 16, 59, 60].
+Our model has several limitations that should be addressed in future work.
+First, it focuses exclusively on algorithmic mediation, i.e.
+the redistribution of existing cultural traits and does not account for algorithmic innovation, where intelligent systems actively generate new solutions or artifacts.
+Previous research has found strong support for the ability of machines to also take part in the innovation process [61, 62], which future models of cultural evolution should take into account.
+Second, we assume that algorithmic mediation remains static, whereas the behavior of the real-world algorithms co-evolves with human behavior driven by user feedback and iterative improvement of machines [60].
+Third, our model does not include biases in algorithmic mediation, such as leading to unintended distortions in content distribution, all of which are prevalent in real systems [63].
+Furthermore, extending our framework to account for more complex cultural spaces will be an essential next step for capturing the richness of real-world cultural dynamics, including more complex innovation processes such as recombination [39].
+Ultimately, the predictions of our model on the interaction between algorithmic mediation and social learning must stand up to empirical testing through controlled laboratory experiments and observational studies in diverse online and offline environments.
+Authors’ Contributions
+A.C.: conceptualization, investigation, software, visualization, writing—original draft, writing—review and editing; F.B.: conceptualization, investigation, visualization, writing—original draft, writing—review and editing; I.R.: conceptualization, investigation, writing—original draft, writing—review and editing.
+All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
+Acknowledgments
+The authors would like to thank the reviewers for their constructive comments.
+Data Accessibility
+Simulation code to reproduce the results in this paper is publicly available in GitHub repository [64] and permanently archived on Zenodo [65].
+Electronic supplementary material is available online [66].
+Funding Statement
+No additional funding was received for this study.